# Bridging Model Reference Adaptive Control and Data Informativity

### v2

@@ 4.2 Relations between the two data informativity conditions
implying that
$\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}=\begin{bmatrix}X_{-}(t)\\ X_{+}(t)\end{bmatrix}\begin{bmatrix}V_{1}&V_{2}\end{bmatrix}$.
Thus, (11) holds. Then, according to Lemma 3, $(U_{-}(t),X(t))$ are informative for model reference control.
Proposition 4.2 shows that, when a solution of the matching equations exists, the full rank condition (4) is a sufficient condition for informativity for model reference control. The following theorem provides conditions under which the full rank condition (4) becomes a necessary condition for informativity for model reference control
@@
Theorem 4.4 establishes the special cases when unique system identification becomes necessary to solve the data-driven model reference control problem, which can be verified by simply checking the size and the rank of $B_{\rm m}$. We refer the reader to the discussion in Section 5.4 on how Theorem 4.4 offers insightful views on existing solutions to the model reference adaptive control problem.
@@ 5 MRAC with Online Informative Data
We aim to design an adaptive law to guarantee asymptotic convergence of $\hat{K}$ and $\hat{L}$ in (7) to some (possibly non-unique) solution of the matching equations (6). We first provide necessary and sufficient conditions for existence of a time at which the system data become informative in the sense of Lemma 3 (cf. Section V-A). We then provide necessary and sufficient conditions for asymptotic convergence to a solution of the matching equations by only relying on data informativity for model reference control (cf. Section V-B). Finally, stability and convergence for the closed-loop signals are studied (cf. Section V-C) and discussions on existing MRAC solutions are given (cf. Section V-D).
5.1 Necessary and sufficient conditions for existence of informative time
To achieve online adaptive control in the data informativity framework, it is important to find the first time instant $t$ at which the data $(U_{-}(t),X(t))\in\mathfrak{B}_{t}$ are informative for model reference control. This motivates the following definition.
@@ 5.1 Necessary and sufficient conditions for existence of informative time
We will now show that if the trajectory $(u,x)\in\mathfrak{B}$ is generated online via an appropriate estimation and control mechanism, then the associated $T^{*}$ is finite if and only if the matching equations (6) have a solution. The estimation and control mechanism will make use of the data matrices
@@
where $\sigma\!\in\!\mathbb{R}$ is a given upper bound on the norm of the collected state data. The use of data matrices with past and current data is inspired by the concurrent learning approach [chowdhary2010concurrent], where past and current data are concurrently used to update the control gains. Note that to construct the data matrices in (20) and (21), we need to verify at every time $t$ whether $t\leq T^{*}+1$. We emphasize that, even though the exact value of $T^{*}$ may be unknown at time $t$, one can verify whether $t\leq T^{*}+1$ via the linear algebraic condition (11).
@@
With the adaptive law in (23), we propose the following adaptive controller:
$u_{\mathrm{a}}(t)=\Phi_{U_{-}}(t)\Theta(t)\begin{bmatrix}x(t)\\ r(t)\end{bmatrix}$. (25)
The idea is to apply a control input that switches between the adaptive controller $u_{\mathrm{a}}(t)$ in (25) and an input $u_{\rm r}(t)$ that increases the rank of the data matrix. We introduce the control input formally in the next lemma, where we also clarify the relation between $T^{*}$ and the existence of solutions of the matching equations (6).
@@
    • ii)
      $\begin{bmatrix}x(t)\\ u_{\mathrm{a}}(t)\end{bmatrix}\in\operatorname{im}\begin{bmatrix}\Phi_{X_{-}}(t)\\ \Phi_{U_{-}}(t)\end{bmatrix}$, and
@@
    then there exists a $u_{\mathrm{r}}(t)\in\mathbb{R}^{m}$ such that
    $\operatorname{rank}\left(\begin{bmatrix}\Phi_{X_{-}}(t)&x(t)\\ \Phi_{U_{-}}(t)&u_{\mathrm{r}}(t)\end{bmatrix}\right)=\operatorname{rank}\left(\begin{bmatrix}\Phi_{X_{-}}(t)\\ \Phi_{U_{-}}(t)\end{bmatrix}\right)+1.$ (26)
    In this case, choose $u(t)=u_{\mathrm{r}}(t)$.
  • Otherwise, choose $u(t)=u_{\mathrm{a}}(t)$ with $u_{\mathrm{a}}(t)$ as in (25).
Consider the resulting input-state trajectory $(u,x)\in\mathfrak{B}$. Then, $T^{*}$ is finite if and only if (6) holds for some $(K,L)$. In this case, $T^{*}\leq n+m$.
Proof 5.8.
We first show that if i), ii) and iii) hold, then there exists a $u_{\mathrm{r}}(t)\in\mathbb{R}^{m}$ satisfying (26). To do so, note that iii) implies that $t\leq T^{*}+1$ so $\Phi_{U_{-}}(t)=U_{-}(t)$ and $\Phi_{X_{-}}(t)=X_{-}(t)$. Moreover, ii) implies that $x(t)\in\operatorname{im}X_{-}(t)$. Therefore, it follows from [van2021beyond, Theorem 1] that there exist a $\xi\in\mathbb{R}^{n}$ and a nonzero $\eta\in\mathbb{R}^{m}$ such that
Hence, we can choose $u_{\mathrm{r}}(t)$ such that
$\xi^{\top}x(t)+\eta^{\top}u_{\mathrm{r}}(t)\neq 0,$ (28)
so that (26) holds. Now we show that $T^{*}$ is finite if and only if (6) has a solution $(K,L)$.
Sufficiency:
Assume that there exists a solution $(K,L)$ of the matching equations (6). We claim that $T^{*}\leq n+m$. Suppose on the contrary that $T^{*}>n+m$. We will prove that
@@
Now, if
$\begin{bmatrix}x(t^{\prime})\\ u_{\mathrm{a}}(t^{\prime})\end{bmatrix}\in\text{im}\begin{bmatrix}\Phi_{X_{-}}(t^{\prime})\\ \Phi_{U_{-}}(t^{\prime})\end{bmatrix}$, (31)
then (30) follows from (26). Otherwise, if the inclusion (31) does not hold, then (30) immediately follows from the fact that the input $u(t^{\prime})$ is chosen as $u(t^{\prime})=u_{a}(t^{\prime})$. This proves that (29) holds. However, by Lemma 2 and Proposition 4.2, we see that $(U_{-}(n+m),X_{-}(n+m))$ are informative for model reference control. This contradicts the definition of $T^{*}$. Therefore, $T^{*}\leq n+m$.
Necessity:
Suppose that $T^{*}$ is finite. Then, according to Definition 5.6 and Lemma 3, there exist gains $(K,L)$ such that $\Sigma_{(U_{-}(T^{*}),X(T^{*}))}\subseteq\Sigma^{K,L}$. Since $(A_{\rm s},B_{\rm s})\in\Sigma_{(U_{-}(T^{*}),X(T^{*}))}$, we have that (6) holds.
@@
(Design of the input). Following [van2021beyond], the input $u_{r}(t)$ in Lemma 5.7 can be designed as follows: first, find $\xi\in\mathbb{R}^{n}$, $\eta\in\mathbb{R}^{m}\setminus\{0\}$ such that (27) holds; then, design $u_{\mathrm{r}}(t)$ such that (28) holds.
@@
(Design of the parameters). In (20) and (21), $\sigma$ can be designed, for instance, as an upper bound of $x_{\rm m}$, which can be determined since the reference model $A_{\rm m},B_{\rm m}$ and the reference input $r\in\ell_{\infty}^{p}$ are given. In (23), the gain has been taken as a scalar $\gamma\in(0,2)$, for simplicity [^2] Requirements analogous to $\gamma\in(0,2)$ are common in discrete-time gradient-based adaptation [ioannou2006adaptive, Lemma 4.1.1]. It will be shown in the proof of Theorem 5.12 that this choice of $\gamma$ also ensures the existence of a limit of the Lyapunov function, which in turn guarantees (10). . However, the scalar $\gamma$ can be generalized to a matrix gain $\Gamma=\Gamma^{\top}$ satisfying $0<\Gamma<2I$, which allows different adaptation rates along different directions [ioannou2006adaptive].
5.2 Necessary and sufficient conditions for convergence to a matching solution
We are now in a position to address the MRAC problem as formulated in Section 3.
@@ 5.2 Necessary and sufficient conditions for convergence to a matching solution
Let $\Phi_{U_{-}}(t)$ and $\Phi_{X}(t)$ be as in (20) and (21), and consider the adaptive law (23) and the input $u(t)$ in Lemma 5.7. Then, the control gains
$\begin{bmatrix}\hat{K}(t)&\hat{L}(t)\end{bmatrix}=\Phi_{U_{-}}(t)\Theta(t)$ (33)
are such that (10) holds if and only if $T^{*}$ is finite.
Proof 5.13.
Sufficiency:
Suppose that $T^{*}$ is finite. By definition of $\Phi_{X}(t)$, there exists a matrix $\Theta^{*}\in\mathbb{R}^{(T^{*}+1)\times(n+p)}$ such that
@@
Necessity:
Suppose that (10) holds. Then,
Since for all $t\in\mathbb{N}$,
and $\operatorname{im}B_{\rm s}$ is closed, we have that
Therefore, there exist gains $(K,L)$ such that $A_{\rm m}-A_{\rm s}=B_{\rm s}K$ and $B_{\rm m}=B_{\rm s}L$, that is, (6) holds. According to Lemma 5.7, $T^{*}$ is finite.
Existing adaptive control methods with relaxed requirements imposed on data like [lavretsky2009combined, chowdhary2010concurrent, chowdhary2013concurrent, cho2017composite, lee2019concurrent, roy2017combined, katiyar2022initial] identify conditions weaker than PE: yet, only sufficiency of such conditions is proven. The key challenge solved by Theorem 5.12 is to provide a necessary and sufficient condition. The sufficiency part of the proof of Theorem 5.12 shows that
$\operatorname{im}\begin{bmatrix}I&0\\ A_{\rm m}&B_{\rm m}\end{bmatrix}\subseteq\operatorname{im}\begin{bmatrix}X_{-}(T^{*})\\ X_{+}(T^{*})\end{bmatrix}$, (43)
implies (34) and thus the asymptotic convergence in (10). Then, the proof of Theorem 5.12 also shows the necessity of (43), where Lemma 5.7 serves to connect the notion of informative data with the matching equations.
Hence, Theorem 5.12 establishes—for the first time—a necessary and sufficient condition for the asymptotic convergence of the adaptive gains to a solution of the matching equations. Notably, the proposed MRAC approach requires neither unique system identification nor structural assumptions on the system apart from controllability. We refer the reader to Section V-D for a discussion on structural assumptions in existing MRAC approaches, such as square assumptions on $L$ ($p=m$), or $B_{{\rm s}}$ being known, or $B_{{\rm s}}$ being full column rank. Before this, we discuss the boundedness of the closed-loop signals and the convergence of the error.
@@ 5.3 Stability and boundedness of the closed-loop signals
([zhou2017asymptotic]). Consider a linear time-varying autonomous system
@@
([jiang2001input]). Consider a nonlinear system
@@
Since $\lim_{t\to\infty}A_{\rm s}+B_{\rm s}\hat{K}(t)=A_{\rm m}$ and $A_{\rm m}$ is Schur, [zhou2017asymptotic, Corollary 3] implies that (47) is uniformly exponentially stable. From Definition 5.14, we have that for any $t_{0}\in\mathbb{Z}_{+}$ and $t>t_{0}$, there exist scalars $\rho^{\prime}>0$ and $\lambda^{\prime}\in(0,1)$ such that
@@
Finally, we demonstrate the convergence of the error. According to [jiang2001input, Sect. 3.2], for the input-to-state stable system (49), there exists a $\mathcal{K}$-function $f^{\prime}$ such that for any $e(0)\in\mathbb{R}^{n}$,
@@ 5.4 Discussion on excitation and structural assumptions in state-of-the-art MRAC
We now discuss the requirements on collected data imposed by the proposed MRAC method as compared to existing state-of-the-art MRAC methods. Without going into full details, we mention that state-of-the-art approaches like concurrent learning or composite MRAC [lavretsky2009combined, chowdhary2010concurrent, chowdhary2013concurrent, cho2017composite, lee2019concurrent, roy2017combined] are based on a Lyapunov design in continuous time, based on the error dynamics
@@
with $A_{\rm m}$ the Hurwitz state matrix of the reference model, $B_{\rm s}$ the control matrix of the system, $\psi$ measured signals, and $\tilde{\Lambda}$ a parameter estimation error. In this paper, we have proposed a discrete-time MRAC method. Although a direct comparison to the state-of-the-art is thus not possible, some considerations can still be made based on the analogy between the error dynamics in (49) and (51). To this end, note that all state-of-the-art solutions with the exception of [roy2017combined] require the input matrix $B_{\rm s}$ to be known, which is not required in the MRAC method in the current paper. The estimation of $B_{\rm s}$ in [roy2017combined] leveraged the concept of initial excitation to relax the classical persistence of excitation conditions. When written in discrete time, initial excitation requires the existence of $T\in\mathbb{N}$ and $\delta>0$ such that
@@
satisfy the full rank condition (4). Thus, from Lemma 2, we conclude that the requirements imposed on collected data by initial excitation are such that $(A_{\rm s},B_{\rm s})$ can be uniquely identified. Such an observation further illustrates the significance of Proposition 4.2 and Theorem 4.4. Similar to [roy2017combined], let us assume that the matching equations (6) have a solution and $p=m$. From Proposition 4.2, we know that (52) implies data informativity for model reference control. However, when $B_{\rm m}$ does not have full column rank, Theorem 4.4 implies that the initial excitation (52) is sufficient but not necessary to solve the model reference control problem. This means that initial excitation imposes stronger conditions on collected data, not required in the proposed MRAC approach.
Let us finally note that most approaches [tao2014multivariable, lavretsky2009combined, chowdhary2010concurrent, chowdhary2013concurrent, cho2017composite, roy2017combined, lee2019concurrent, katiyar2022initial] in the multi-variable MRAC literature require $B_{\rm s}$ to have full column rank, which in turns guarantees a unique solution $(K,L)$ of the matching equations (6). This is not required in the proposed MRAC method, which remains valid even when an infinite number of solutions of the matching equations (6) exist. This will be illustrated in the numerical example in the next section.

### v3

@@ 4.2 Relations between the two data informativity conditions
implying (12), that is, (11) holds. Then, according to Lemma 3, $(U_{-}(t),X(t))$ are informative for model reference control.
Proposition 4.2 shows that, when a solution of the matching equations exists, the full rank condition (4) is a sufficient condition for informativity for model reference control. The following theorem provides conditions under which the full rank condition (4) becomes a necessary condition for informativity for model reference control.
@@
Theorem 4.4 establishes the special cases when unique system identification becomes necessary to solve the data-driven model reference control problem, which can be verified by simply checking the size and rank of $B_{\rm m}$. We refer the reader to the discussion in Section 5.5 on how Theorem 4.4 offers insightful views on existing solutions to the MRAC problem.
@@ 5 MRAC with Online Informative Data
To solve Problem 1, we first present necessary and sufficient conditions for the existence of an adaptive law that guarantees asymptotic convergence of $\hat{K}(t)$ and $\hat{L}(t)$ in (7) to some (possibly non-unique) solution of the matching equations (6) (Section V-A). We then develop an input design only relying on data informativity for model reference control (Section V-B). Next, based on this input design, we establish convergence of MRAC to a solution of the matching equations (Section V-C). Finally, stability and boundedness for the closed-loop signals are studied (Section V-D) and discussions on existing MRAC solutions are given (Section V-E).
5.1 Necessary and sufficient conditions for existence
We aim to address the first part of Problem 1, that is, establish necessary and sufficient conditions for the existence of an adaptive law and associated adaptive gains such that (10) holds. Existence is analyzed from two perspectives: a data-driven perspective based on the data generated by the system, and a model-based perspective based on the system matrices.
Theorem 5.6.
The following statements are equivalent:
  • (i)
    There exist an adaptive law as in (8) and associated adaptive gains as in (9) such that (10) holds;
  • (ii)
    There exist $t\in\mathbb{N}$ and $(U_{-}(t),X(t))\in\mathfrak{B}_{t}$ such that $(U_{-}(t),X(t))$ are informative for model reference control;
  • (iii)
    There exist $K\in\mathbb{R}^{m\times n}$ and $L\in\mathbb{R}^{m\times p}$ satisfying the matching equations (6).
Proof 5.7.
[…]
Suppose there exist an adaptive law as in (8) and associated adaptive gains as in (9) such that (10) holds. Then,
Since for all $t\in\mathbb{N}$,
and $\operatorname{im}B_{\rm s}$ is closed, we have
[…]
Suppose there exist $t\in\mathbb{N}$ and $(U_{-}(t),X(t))\in\mathfrak{B}_{t}$ such that $(U_{-}(t),X(t))$ are informative for model reference control. Then, there exist $V_{1}$ and $V_{2}$ such that (12) holds. Let $K=U_{-}V_{1}$ and $L=U_{-}V_{2}$, then
$A+BK$ $=AX_{-}V_{1}+BU_{-}V_{1}=X_{+}V_{1}=A_{\rm m},$
$BL$ $=AX_{-}V_{2}+BU_{-}V_{2}=X_{+}V_{2}=B_{\rm m},$
implying statement (iii).
[…]
This implication follows from Theorem 5.14 in Section 5.3.
5.2 Design of the adaptive law and the input
We now provide a recipe for constructing $i_{t},j_{t},\alpha_{t},\beta_{t}$ and $v$ in Problem 1. To achieve online adaptive control in the data informativity framework, it is important to find the first time instant $t$ at which the data $(U_{-}(t),X(t))\in\mathfrak{B}_{t}$ are informative for model reference control. This motivates the following definition.
@@ 5.2 Design of the adaptive law and the input
We will now show that if the trajectory $(u,x)\in\mathfrak{B}$ is generated online via an appropriate estimation and control mechanism, then the associated $T^{*}\leq n+m$ if and only if the matching equations (6) have a solution. The estimation and control mechanism will make use of the data matrices
@@
where $\sigma\!\in\!\mathbb{R}$ is a given upper bound on the norm of the collected state data. The use of data matrices with past and current data is inspired by the concurrent learning approach [8], where past and current data are concurrently used to update the control gains. Note that to construct the data matrices in (21) and (22), we need to verify at every time $t$ whether $t\leq T^{*}+1$. We emphasize that, even though the exact value of $T^{*}$ may be unknown at time $t$, one can verify whether $t\leq T^{*}+1$ via the linear algebraic condition (11).
@@
Given the adaptive law in (24), we introduce the control input formally in the next lemma, where we also clarify the relation between $T^{*}$ and the existence of solutions of the matching equations (6).
@@
[…]
@@
    then there exists a $w\in\mathbb{R}^{m}$ such that
    $\operatorname{rank}\left(\begin{bmatrix}\Phi_{X_{-}}(t)&x(t)\\ \Phi_{U_{-}}(t)&w\end{bmatrix}\right)=\operatorname{rank}\left(\begin{bmatrix}\Phi_{X_{-}}(t)\\ \Phi_{U_{-}}(t)\end{bmatrix}\right)+1.$ (27)
    In this case, select $u(t)$ as in (7) with
    $\hat{K}(t)=0,\hat{L}(t)=0\text{ and }v(t)=w.$ (28)
  • Otherwise, select $u(t)$ as in (7) with
    $\begin{bmatrix}\hat{K}(t)&\hat{L}(t)\end{bmatrix}=\Phi_{U_{-}}(t)\Theta(t)\text{ and }v(t)=0.$ (29)
Consider the resulting input-state trajectory $(u,x)\in\mathfrak{B}$. Then, $T^{*}\leq n+m$ if and only if (6) holds for some $(K,L)$.
Proof 5.10.
We first show that if i), ii) and iii) hold, then there exists a $w\in\mathbb{R}^{m}$ satisfying (27). To do so, note that iii) implies that $t\leq T^{*}+1$ so $\Phi_{U_{-}}(t)=U_{-}(t)$ and $\Phi_{X_{-}}(t)=X_{-}(t)$. Moreover, ii) implies that $x(t)\in\operatorname{im}X_{-}(t)$. Therefore, it follows from [38, Theorem 1] that there exist a $w$ such that (27) holds.
Now we show that $T^{*}\leq n+m$ if and only if condition (iii) in Theorem 5.6 holds.
For necessity, we suppose that $T^{*}\leq n+m$. According to Definition 5.8 and Lemma 3, there exist gains $(K,L)$ such that $\Sigma_{(U_{-}(T^{*}),X(T^{*}))}\subseteq\Sigma^{K,L}$. Since $(A_{\rm s},B_{\rm s})\in\Sigma_{(U_{-}(T^{*}),X(T^{*}))}$, we have that (6) holds.
For sufficiency, we suppose on the contrary that $T^{*}>n+m$. We will prove that
@@
Now, if (26) holds for $t=t^{\prime}$, then (31) follows from selecting $u(t^{\prime})$ as in (28) for $t=t^{\prime}$. Otherwise, if (26) does not hold for $t=t^{\prime}$, then (31) follows from selecting $u(t^{\prime})$ as in (29) for $t=t^{\prime}$. This proves that (30) holds. However, by Lemma 2 and Proposition 4.2, we see that $(U_{-}(n+m),X_{-}(n+m))$ are informative for model reference control. This contradicts the definition of $T^{*}$. Therefore, $T^{*}\leq n+m$.
@@
(Design of the input). Following the approach in [38, Theorem 1], one way to design $w$ is as follows. First, find $\xi\in\mathbb{R}^{n}$, $\eta\in\mathbb{R}^{m}\setminus\{0\}$ such that
Then, choose $w$ such that
$\xi^{\top}x(t)+\eta^{\top}w\neq 0.$ (34)
@@
(Design of the parameters). In (21) and (22), $\sigma$ can be designed, for instance, as an upper bound of $x_{\rm m}$, which can be determined since the reference model $A_{\rm m},B_{\rm m}$ and the reference input $r\in\ell_{\infty}^{p}$ are given. In (24), the gain has been taken as a scalar $\gamma\in(0,2)$, for simplicity [^2] Requirements analogous to $\gamma\in(0,2)$ are common in discrete-time gradient-based adaptation [19, Lemma 4.1.1]. It will be shown in the proof of Theorem 5.14 that this choice of $\gamma$ also ensures the existence of a limit of the Lyapunov function, which in turn guarantees (10). . However, the scalar $\gamma$ can be generalized to a matrix gain $\Gamma=\Gamma^{\top}$ satisfying $0<\Gamma<2I$, which allows different adaptation rates along different directions [19].
5.3 Convergence to a matching solution
We are now in a position to formulate the convergence properties for the proposed recipe.
@@ 5.3 Convergence to a matching solution
If the matching equations (6) have a solution, the adaptive gains designed as in (28)-(29) are such that (10) holds.
Proof 5.15.
By Lemma 5.9, we have $T^{*}\leq n+m$. Then, by definition of $\Phi_{X}(t)$, there exists a matrix $\Theta^{*}\in\mathbb{R}^{(T^{*}+1)\times(n+p)}$ such that
@@
Theorem 5.14 shows that, provided the matching equations (6) admit a solution, the input design in Lemma 5.9 guarantees convergence to a solution of the matching equation by only exploiting data that are informative for model reference control. Notably, the proposed MRAC approach requires neither unique system identification nor structural assumptions on the system apart from controllability. We refer the reader to Section 5.5 for a discussion on structural assumptions in existing MRAC approaches, such as square assumptions on $L$ ($p=m$), or $B_{{\rm s}}$ being known, or $B_{{\rm s}}$ being full column rank. Before this, we discuss the boundedness of the closed-loop signals and the convergence of the error.
@@ 5.4 Stability and boundedness of the closed-loop signals
([47]). Consider a linear time-varying autonomous system
@@
([20]). Consider a nonlinear system
@@
Since $\lim_{t\to\infty}A_{\rm s}+B_{\rm s}\hat{K}(t)=A_{\rm m}$ and $A_{\rm m}$ is Schur, [47, Corollary 3] implies that (47) is uniformly exponentially stable. From Definition 5.16, we have that for any $t_{0}\in\mathbb{Z}_{+}$ and $t>t_{0}$, there exist scalars $\rho^{\prime}>0$ and $\lambda^{\prime}\in(0,1)$ such that
@@
Finally, we demonstrate the convergence of the error. According to [20, Sect. 3.2], for the input-to-state stable system (49), there exists a $\mathcal{K}$-function $f^{\prime}$ such that for any $e(0)\in\mathbb{R}^{n}$,
@@ 5.5 Discussion on excitation and structural assumptions in state-of-the-art MRAC
We now discuss the requirements on collected data imposed by the proposed MRAC method as compared to existing state-of-the-art MRAC methods. Without going into full details, we mention that state-of-the-art approaches like concurrent learning or composite MRAC [25, 8, 9, 7, 26, 31] are based on a Lyapunov design in continuous time, based on the error dynamics
@@
with $A_{\rm m}$ the Hurwitz state matrix of the reference model, $B_{\rm s}$ the control matrix of the system, $\psi$ measured signals, and $\tilde{\Lambda}$ a parameter estimation error. In this paper, we have proposed a discrete-time MRAC method. Although a direct comparison to the state-of-the-art is thus not possible, some considerations can still be made based on the analogy between the error dynamics in (49) and (51). To this end, note that all state-of-the-art solutions with the exception of [31] require the input matrix $B_{\rm s}$ to be known, which is not required in the MRAC method in the current paper. The estimation of $B_{\rm s}$ in [31] leveraged the concept of initial excitation to relax the classical persistence of excitation conditions. When written in discrete time, initial excitation requires the existence of $T\in\mathbb{N}$ and $\delta>0$ such that
@@
satisfy the full rank condition (4). Thus, from Lemma 2, we conclude that the requirements imposed on collected data by initial excitation are such that $(A_{\rm s},B_{\rm s})$ can be uniquely identified. Such an observation further illustrates the significance of Proposition 4.2 and Theorem 4.4. Similar to [31], let us assume that the matching equations (6) have a solution and $p=m$. From Proposition 4.2, we know that (52) implies data informativity for model reference control. However, when $B_{\rm m}$ does not have full column rank, Theorem 4.4 implies that the initial excitation (52) is sufficient but not necessary to solve the model reference control problem. This means that initial excitation imposes stronger conditions on collected data, not required in the proposed MRAC approach. Indeed, adaptive approaches based on initial excitation or other relaxed excitation conditions [25, 8, 9, 7, 26, 31] only prove the sufficiency of such conditions for converging to a solution of the matching equations. In comparison to these works, the contribution of this paper is to establish necessary and sufficient conditions for converging to a solution of the matching equations.
The works [24, 23] study least squares estimation of the system parameters in the absence of persistence of excitation. These works show that, in this case, the parameters do not necessarily converge to the true system parameters. However, the closed-loop system matrix may converge to a reference matrix. The latter convergence, however, is not always guaranteed and depends on the quality of the data, which was not discussed in [24, 23]. In comparison to these works, the contribution of this paper is to highlight the required conditions on the closed-loop data under which the adaptive gains converge to a solution of the matching equations.
Let us finally note that most approaches [25, 8, 9, 7, 31, 26] in the multi-variable MRAC literature require $B_{\rm s}$ to have full column rank, which in turns guarantees a unique solution $(K,L)$ of the matching equations (6) [35]. This is not required in the proposed MRAC method, which remains valid even when an infinite number of solutions of the matching equations (6) exist. This will be illustrated in the numerical example in the next section.